# The stable trees revisited

### v1

@@ 5 Relation to Wang’s construction
In this section, we show that our construction provides a new perspective on Wang’s construction of the stable tree from [minmin]. The key result is Proposition 5.4 below, which connects $\widetilde{\sigma}$ and the jumps of a normalised $\alpha$-stable excursion.
For a sequence $\mathbf{x}=(x_{1},x_{2},\ldots)$ with a largest element, we will write $(\mathbf{x})^{\downarrow}$ for the same sequence with the entries put into decreasing order. Recall that $\Delta_{1}\geq\Delta_{2}\geq\dots\geq 0$ are the ordered jumps of $\widetilde{\sigma}$. Let $\Delta_{1}^{n}\geq\Delta_{2}^{n}\geq\dots\geq 0$ denote the ordered (out-)degrees in the finite-$n$ model, so that $(\Delta_{1}^{n},\Delta_{2}^{n},\ldots,\Delta_{n}^{n})=(D_{1},D_{2},\ldots,D_{n})^{\downarrow}$. For all $m\geq n+1$, let $\Delta^{n}_{m}=0$ so that we append an infinite sequence of 0’s to the end of this vector.
Let $\ell_{2}^{\downarrow}=\{(x_{1},x_{2},\ldots):x_{1}\geq x_{2}\ldots\geq 0\}$ which we endow with the usual $\ell_{2}$ norm. For sequences indexed by an arbitrary countable set $\Gamma$ we instead write $\ell_{2}(\Gamma)$ for the corresponding set of square summable sequences, and $\ell_{2}^{+}(\Gamma)$ when all of the elements of the sequences are strictly positive.
@@
The proof of this result will make use of the theory of size-biased point processes developed by Aldous in Section 3.3 of [aldouscritrg]. We need a little set-up.
@@
Observe we may also construct the process $S(t)$ in terms of an i.i.d. sequence of $\textnormal{Exp}\left(1\right)$ random variables, say $\{E_{\gamma}^{\prime}:\gamma\in\Gamma\}$, again independent of $\mathbf{Y}$: then we may set
@@
A convenient formulation of the result we shall use is given in Proposition 17 of Aldous and Limić [aldouslimic], which we reproduce below.
@@
To prove this, by a strengthening of Theorem 16.16 of [kallenbergfmp] it is enough to prove convergence in distribution of $\Xi^{(n)}(U)$ to $\Xi^{(\infty)}(U)$ for any open, relatively compact set $U$ such that $\Xi^{(\infty)}(\partial U)=0$ a.s. Proving this statement is routine but rather long, so we give only a brief sketch here: fix a set $U$ and let us work on a common probability space such that the convergence $(\frac{1}{a_{n}}\sum_{i=1}^{\lfloor tm_{n}\rfloor}\widehat{D}_{i},t\geq 0)\to\widetilde{\sigma}$ in $D([0,\infty))$ is almost sure. We now work on the (probability 1) event that this convergence does occur and that the limiting point measure puts no mass on the boundary.
@@
It remains to show that for all but finitely many $n$ these are the only atoms in $U$. Suppose not, then we can find a subsequence in $n$ along which there is an extra atom $(u^{(n)},v^{(n)})\in U$. Using relative compactness of $U$ it is straightforward to derive a contradiction: either we find another atom in the limiting measure in $U$ or on the boundary, or we find that $\Xi^{(n)}$ has a pair of atoms that both converge to a common point (which violates either right continuity or existence of left limits in $\widetilde{\sigma}$).
@@
converge to $\rho$ in probability on the space $D([0,T])$. Moreover, the jump chain of $\rho^{n}$ has the same distribution as that of $\widetilde{\sigma}^{n}$. Let $0=\tau_{0}^{n}<\tau_{1}^{n}<\tau_{2}^{n}<\dots<\tau_{n}^{n}$ be the jump times of $\rho^{n}$, and let $(\widehat{\Delta}^{n})$ be the reordering of $(\Delta^{n})$ induced by the jumps of $\rho^{n}$. We see that $\widetilde{\sigma}^{n}(t)\overset{d}{=}\rho^{n}(\lambda^{n}(t))$, where $\lambda^{n}$ is the unique piecewise linear function with $\lambda^{n}(\frac{i}{m_{n}})=\tau_{i}$, and (say) such that $\lambda^{n}$ is constant beyond time $n/m_{n}$.
We would like to use $\lambda^{n}$ as a time change for bounding the Skorokhod distance between $\rho^{n}$ and $\sigma^{n}$, but unfortunately it may not fix the endpoint $T$. In order to resolve this issue, we define a modified version $\widetilde{\lambda}^{n}$ such that
$\widetilde{\lambda}^{n}\left(\frac{i}{m_{n}}\right)=\tau_{i}\wedge T\quad\text{ for }i\leq\lfloor Tm_{n}\rfloor$
@@
$\sup_{t\in[0,T]}|\widetilde{\lambda}^{n}(t)-t|\overset{p}{\to}0\quad\text{ and }\quad\sup_{t\in[0,T]}|\rho^{n}\circ\lambda^{n}-\rho^{n}\circ\widetilde{\lambda}^{n}|\overset{p}{\to}0,$
which will establish $\rho^{n}\overset{p}{\to}\widetilde{\sigma}$ in the Skorokhod metric. For the first limit, we note that we can bound the supremum by looking only at the endpoints of each piece. Moreover, for all $i\leq\lfloor Tm_{n}\rfloor$ note that $\frac{i}{m_{n}}$ is closer to $\tau_{i}\wedge T$ than to $\tau_{i}$. Hence we can bound
$\sup_{t\in[0,T]}|\widetilde{\lambda}^{n}(t)-t|\leq\sup_{i\leq\lfloor Tm_{n}\rfloor}\left|\tau_{i}-\frac{i}{m_{n}}\right|.$
By identifying the conditional law of $\tau_{i}-\tau_{i-1}$, we see that the process defined by
$Y_{i}=\tau_{i}-\frac{1}{m_{n}}\sum_{j=1}^{i}\left(1-\frac{1}{n}-\frac{1}{n}\sum_{k=1}^{j-1}\widehat{\Delta}_{k}^{n}\right)^{-1}$
is a mean 0 martingale. We may bound
$\mathbb{P}\left(\sup_{i}\left|\tau_{i}-\frac{i}{m_{n}}\right|>\varepsilon\right)\\ \leq\mathbb{P}\left(\sup_{i}|Y_{i}|>\frac{1}{2}\varepsilon\right)+\mathbb{P}\left(\sup_{i}\frac{1}{m_{n}}\left|\sum_{j=1}^{i}\left(1-\frac{1}{n}-\frac{1}{n}\sum_{k=1}^{j-1}\widehat{\Delta}_{k}^{n}\right)^{-1}-i\right|>\frac{1}{2}\varepsilon\right).$ (14)
@@
$\mathbb{P}\left(\left\{\sup_{i}|Y_{i}|>\frac{1}{2}\varepsilon\right\}\cap A_{n}\right)\leq\frac{16}{\varepsilon^{2}}\mathbb{E}\left[Y_{\lfloor Tm_{n}\rfloor}^{2}\mathbbm{1}\left(A_{n}\right)\right]\leq\frac{16}{\varepsilon^{2}}\sum_{i=1}^{\lfloor Tm_{n}\rfloor}\mathbb{E}\left[(Y_{i}-Y_{i-1})^{2}\mathbbm{1}\left(A_{n}\right)\right].$
Now for each $i$ we can estimate
@@
It remains to bound the distance between $\rho^{n}(\lambda^{n}(t))$ and $\rho^{n}(\widetilde{\lambda}^{n}(t))$. Observe that this distance is maximised at $t=T$ (we prove this by splitting into the cases $\tau_{\lfloor Tm_{n}\rfloor}\leq T$ and $\tau_{\lfloor Tm_{n}\rfloor}>T$), where we have
$\rho^{n}(\lambda^{n}(T))=\frac{1}{m_{n}}\sum_{i=1}^{\lfloor Tm_{n}\rfloor}\widehat{\Delta}_{i}^{n}\quad\text{ and }\rho^{n}(\widetilde{\lambda}^{n}(T))=\frac{1}{m_{n}}\sum_{i=1}^{I}\widehat{\Delta}_{i}^{n},$
where $I=\sup\{i:\tau_{i}\leq T\}$. For any $\delta>0$ the event $B_{\delta}^{n}=\{I\in((T-\delta)m_{n},(T+\delta)m_{n})\}$ has probability tending to 1, and on this event we have
$|\rho^{n}(\widetilde{\lambda}^{n}(T))-\rho^{n}(\lambda^{n}(T))|\leq\frac{1}{m_{n}}\sum_{i=\lfloor(T-\delta)m_{n}\rfloor}^{\lfloor(T+\delta)m_{n}\rfloor}\widehat{\Delta}_{i}^{n}.$
As $n\to\infty$ the right-hand side converges in distribution to $\widetilde{\sigma}_{T+\delta}-\widetilde{\sigma}_{T-\delta}$. Thus, by picking $\delta$ suitably for each $\varepsilon$, we see that $\mathbb{P}\left(|\rho^{n}(\widetilde{\lambda}^{n}(T))-\rho^{n}(\lambda^{n}(T))|>\varepsilon\right)\to 0$. This proves $\widetilde{\sigma}^{n}\overset{d}{\to}\rho$.
Thus we have identified $\widetilde{\sigma}_{t}\overset{d}{=}\sum_{i=1}^{\infty}\Delta_{i}\mathbbm{1}\left(E_{i}\leq\Delta_{i}t\right)$. It remains only to identify $(\Delta_{i})$. By 5.1 we know that for any $k$ we have $a_{n}^{-1}(\Delta_{1}^{n},\dots,\Delta_{k}^{n})\overset{d}{\to}(\Delta_{1},\dots,\Delta_{k})$, so we will show the same sequence converges to $(\Theta_{1},\dots,\Theta_{k})$. Let $L^{n}:\{0,1,\dots,n\}\to\mathbb{Z}$ be the Łukasiewicz path of the discrete tree $T^{n}$ of size $n$, and construct a càdlàg function $e^{n}\in D([0,1])$ by taking
@@
Then we have $e^{n}\overset{d}{\to}\mathbbm{e}$, where $\mathbbm{e}$ is a normalised $\alpha$-stable excursion (this result is implicit in [duquesnealphalimit] or [igorduquesneproof]). Now consider a map $\mathfrak{J}:D([0,1])\to\mathbb{R}^{k}$ given as follows: for any $f\in D([0,1])$, we identify the $k$ largest upward jumps of $f$ and arrange them into decreasing order (appending zeros if there are fewer than $k$ such jumps). This gives a well-defined continuous map, and applying it to the above convergence gives
@@
Acknowledgements
We are grateful to Andreas Kyprianou for a very helpful suggestion in the early stages of the project, and to Minmin Wang for a discussion of [minmin]. LH’s work is supported by the EPSRC Centre for Doctoral Training in the Mathematics of Random Systems: Analysis, Modelling, and Simulation (EP/S023925/1). Some of the work was carried out while CG was in residence at the Simons Laufer Mathematical Sciences Institute in Berkeley, California, during the Spring 2025 semester, which LH also visited for a week with funding from the Institute. They gratefully acknowledge the financial support of the National Science Foundation under Grant No. DMS-1928930.
Appendix: omitted proofs
@@ Acknowledgements
First observe that $\sigma$ is killed at rate $a$, independently of its other jumps, and so $\mathbb{P}\left(\sigma_{t}<\infty\right)=e^{-at}$; we can pull this factor out of the expectation and henceforth forget about the killing in $\sigma$: we aim to show that
@@
To this end, set $G(z)=icz-\Psi_{L}(z)$, $z\in\mathbb{C}$. We can show that, for any closed polygonal contour $\gamma$, we have $\oint_{\gamma}G(z)\,dz=0$ — this is by Fubini’s theorem, which is valid due to condition (i). Thus by Morera’s theorem $G$ is holomorphic on $\mathbb{C}$, so $\exp(G(z))$ is holomorphic too.
@@
To this end, we will show that the integrands of these two integrals tend to 0 uniformly: that is, for $z=\lambda-is$, $s\in[0,t]$ we have $\exp(G(z))\to 0$ uniformly in $s$ as $|\lambda|\to\infty$. As $|\exp(G(z))|=\exp(\textnormal{Re}\left(G(z)\right))$ it suffices to show $\textnormal{Re}\left(G(z)\right)\to$-$\infty$ uniformly. It is straightforward to calculate
@@
$\mathbb{E}\left[\left|\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}T\right)+\mathbbm{1}\left(E_{i}\leq x_{i}T\right)\right)\right|\right]$ $\leq\sum_{i=N+1}^{\infty}x_{i}\left(\mathbb{P}\left(E_{i}\leq x_{i}^{(n)}T\right)+\mathbb{P}\left(E_{i}\leq x_{i}T\right)\right)$
$\leq T\sum_{i=N+1}^{\infty}x_{i}(x_{i}^{(n)}+x_{i})$
@@
which converges to 0 almost surely (on the event that $\lambda^{(n)}$ is defined). Letting $A_{3}^{(n)}$ be the event that $\sup_{t\in[0,T]}\left|\lambda^{(n)}(t)-t\right|\leq\varepsilon$, we see that $\mathbb{P}\left(A_{3}^{(n)}\right)\to 1$. Note that $\widetilde{R}^{(n)}(\lambda^{(n)}(t))=\widetilde{S}(t)$ for all $t\in[0,T]$. In particular, we can compute
$\sup_{t\in[0,T]}|R^{(n)}(\lambda^{(n)}(t))-S(t)|$ $=\sup_{t\in[0,T]}\left|\sum_{i=1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}t\right)-\mathbbm{1}\left(E_{i}\leq x_{i}t\right)\right)\right|$
$=\sup_{t\in[0,T]}\left|\widetilde{R}^{(n)}(t)-\widetilde{S}(t)+\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}t\right)-\mathbbm{1}\left(E_{i}\leq x_{i}t\right)\right)\right|$
$=\sup_{t\in[0,T]}\left|\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}t\right)-\mathbbm{1}\left(E_{i}\leq x_{i}t\right)\right)\right|$
$\leq\left|\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}T\right)+\mathbbm{1}\left(E_{i}\leq x_{i}T\right)\right)\right|.$
@@
This completes the proof in the case that $\mathbf{X}^{(n)}$ and $\mathbf{X}$ are deterministic. For the general case, we use Skorokhod’s representation theorem to construct $\mathbf{X}^{(n)}$ and $\mathbf{X}$ on the same space such that $\mathbf{X}^{(n)}\to\mathbf{X}$ almost surely. Constructing our exponential random variables also on this space, independent of everything else, and using them to define all processes $S^{(n)}$ and $S$, we deduce from the previous case that for any fixed $\varepsilon>0$ we have

### v2

@@ 5 Relation to Wang’s construction
In this section, we show that our construction provides a new perspective on Wang’s construction of the stable tree from [41]. The key result is 5.4 below, which connects $\widetilde{\sigma}$ and the jumps of a normalised $\alpha$-stable excursion.
Suppose $\mathbf{x}=(x_{1},x_{2},\ldots)$ is a positive real sequence such that, for any $\varepsilon>0$, the set $\{i\in\mathbb{N}:x_{i}\geq\varepsilon\}$ is finite. Then we can rearrange the terms into decreasing order: write $(\mathbf{x})^{\downarrow}$ for this rearrangement. This notion clearly generalises to families $(x_{\gamma}:\gamma\in\Gamma)$ of positive real numbers indexed by arbitrary countable sets. Recall that $\Delta_{1}\geq\Delta_{2}\geq\dots\geq 0$ are the ordered jumps of $\widetilde{\sigma}$. Let $\Delta_{1}^{n}\geq\Delta_{2}^{n}\geq\dots\geq 0$ denote the ordered (out-)degrees in the finite-$n$ model, so that $(\Delta_{1}^{n},\Delta_{2}^{n},\ldots,\Delta_{n}^{n})=(D_{1},D_{2},\ldots,D_{n})^{\downarrow}$. For all $m\geq n+1$, let $\Delta^{n}_{m}=0$ so that we append an infinite sequence of 0’s to the end of this vector.
Let $\ell_{2}^{\downarrow}=\{(x_{1},x_{2},\ldots)\in\ell_{2}:x_{1}\geq x_{2}\ldots\geq 0\}$ which we endow with the usual $\ell_{2}$ norm. For sequences indexed by an arbitrary countable set $\Gamma$ we instead write $\ell_{2}(\Gamma)$ for the corresponding set of square summable sequences, and $\ell_{2}^{+}(\Gamma)$ when all of the elements of the sequences are strictly positive.
@@
The proof of this result will make use of the theory of size-biased point processes developed by Aldous in Section 3.3 of [12]. We need a little set-up.
@@
Observe we may also construct the process $S(t)$ in terms of an i.i.d. sequence of $\textnormal{Exp}\left(1\right)$ random variables, say $\{E_{\gamma}^{\prime}:\gamma\in\Gamma\}$, this time independent of $\mathbf{Y}$: then we may set
@@
A convenient formulation of the result we shall use is given in Proposition 17 of Aldous and Limić [5], which we reproduce below.
@@
To prove this, by a strengthening of Theorem 16.16 of [32] it is enough to prove convergence in distribution of $\Xi^{(n)}(U)$ to $\Xi^{(\infty)}(U)$ for any open, relatively compact set $U$ such that $\Xi^{(\infty)}(\partial U)=0$ a.s. Proving this statement is routine but rather long, so we give only a brief sketch here: fix a set $U$ and let us work on a common probability space such that the convergence $(\frac{1}{a_{n}}\sum_{i=1}^{\lfloor tm_{n}\rfloor}\widehat{D}_{i},t\geq 0)\to\widetilde{\sigma}$ in $D([0,\infty))$ is almost sure. We now work on the (probability 1) event that this convergence does occur and that the limiting point measure puts no mass on the boundary.
@@
It remains to show that for all but finitely many $n$ these are the only atoms of $\Xi^{(n)}$ in $U$. Suppose not, then we can find a subsequence in $n$ along which there is an extra atom $(u^{(n)},v^{(n)})\in U$. Using relative compactness of $U$ it is straightforward to derive a contradiction: either we find another atom in the limiting measure in $U$ or on the boundary, or we find that $\Xi^{(n)}$ has a pair of atoms that both converge to a common point (which violates either right continuity or existence of left limits in $\widetilde{\sigma}$).
@@
converge to $\rho$ in probability on the space $D([0,T])$. Moreover, the jump chain of $\rho^{n}$ has the same distribution as that of $\widetilde{\sigma}^{n}$. Let $0=\tau_{0}^{n}<\tau_{1}^{n}<\tau_{2}^{n}<\dots<\tau_{n}^{n}$ be the jump times of $\rho^{n}$, and let $(\widehat{\Delta}^{n})$ be the reordering of $(\Delta^{n})$ induced by the jumps of $\rho^{n}$. We see that $\widetilde{\sigma}^{n}(t)\overset{d}{=}\rho^{n}(\lambda^{n}(t))$, where $\lambda^{n}$ is the minimal piecewise linear function such that $\lambda^{n}(\frac{i}{m_{n}})=\tau_{i}^{n}$ and (say) $\lambda^{n}$ is constant beyond time $n/m_{n}$. As such, we can explicitly construct $\widetilde{\sigma}^{n}(t):=\rho^{n}(\lambda^{n}(t))$ on our common probability space. We will show that, with this construction, $\widetilde{\sigma}^{n}(t)\overset{p}{\to}\rho$ with respect to the Skorokhod metric.
We aim to bound the Skorokhod distance $d_{D}=d_{D([0,T])}$ between $\widetilde{\sigma}^{n}$ and $\rho^{n}$: let $\widetilde{\lambda}^{n}$ be the minimal piecewise linear time change such that
$\widetilde{\lambda}^{n}\left(\frac{i}{m_{n}}\right)=\tau_{i}^{n}\wedge T\quad\text{ for }i\leq\lfloor Tm_{n}\rfloor$
@@
$\sup_{t\in[0,T]}|\widetilde{\lambda}^{n}(t)-t|\overset{p}{\to}0\quad\text{ and }\quad\sup_{t\in[0,T]}|\widetilde{\sigma}^{n}-\rho^{n}\circ\widetilde{\lambda}^{n}|\overset{p}{\to}0,$
from which we can deduce that $d_{D}(\widetilde{\sigma}^{n},\rho^{n})\overset{p}{\to}0$. For the first limit, we note that we can bound the supremum by looking only at the endpoints of each piece. Moreover, for all $i\leq\lfloor Tm_{n}\rfloor$ note that $\frac{i}{m_{n}}$ is closer to $\tau_{i}^{n}\wedge T$ than to $\tau_{i}^{n}$. Hence we can bound
$\sup_{t\in[0,T]}|\widetilde{\lambda}^{n}(t)-t|\leq\sup_{i\leq\lfloor Tm_{n}\rfloor}\left|\tau_{i}^{n}-\frac{i}{m_{n}}\right|.$
By identifying the conditional law of $\tau_{i}^{n}-\tau_{i-1}^{n}$, we see that the process defined by
$Y_{i}=\tau_{i}^{n}-\frac{1}{m_{n}}\sum_{j=1}^{i}\left(1-\frac{1}{n}-\frac{1}{n}\sum_{k=1}^{j-1}\widehat{\Delta}_{k}^{n}\right)^{-1}$
is a mean 0 martingale, both in its natural filtration $(\mathcal{F}_{i})$ and the larger filtration $\widetilde{\mathcal{F}_{i}}=\sigma(\mathcal{F}_{i},(\widehat{\Delta}_{k}^{n})_{k\geq 1})$. We may bound
$\mathbb{P}\left(\sup_{i}\left|\tau_{i}^{n}-\frac{i}{m_{n}}\right|>\varepsilon\right)\\ \leq\mathbb{P}\left(\sup_{i}|Y_{i}|>\frac{1}{2}\varepsilon\right)+\mathbb{P}\left(\sup_{i}\frac{1}{m_{n}}\left|\sum_{j=1}^{i}\left(1-\frac{1}{n}-\frac{1}{n}\sum_{k=1}^{j-1}\widehat{\Delta}_{k}^{n}\right)^{-1}-i\right|>\frac{1}{2}\varepsilon\right).$ (14)
@@
$\mathbb{P}\left(\left\{\sup_{i}|Y_{i}|>\frac{1}{2}\varepsilon\right\}\cap A_{n}\right)\leq\frac{16}{\varepsilon^{2}}\mathbb{E}\left[Y_{\lfloor Tm_{n}\rfloor}^{2}\mathbbm{1}\left(A_{n}\right)\right]\leq\frac{16}{\varepsilon^{2}}\sum_{i=1}^{\lfloor Tm_{n}\rfloor}\mathbb{E}\left[(Y_{i}-Y_{i-1})^{2}\mathbbm{1}\left(A_{n}\right)\right],$
[…]
@@
It remains to bound the distance between $\rho^{n}(\widetilde{\lambda}^{n}(t))$ and $\widetilde{\sigma}^{n}(t)=\rho^{n}(\lambda^{n}(t))$. Observe that this distance is maximised at $t=T$ (we prove this by splitting into the cases $\tau_{\lfloor Tm_{n}\rfloor}^{n}\leq T$ and $\tau_{\lfloor Tm_{n}\rfloor}^{n}>T$), where we have
$\widetilde{\sigma}^{n}(T)=\frac{1}{m_{n}}\sum_{i=1}^{\lfloor Tm_{n}\rfloor}\widehat{\Delta}_{i}^{n}\quad\text{ and }\rho^{n}(\widetilde{\lambda}^{n}(T))=\frac{1}{m_{n}}\sum_{i=1}^{I}\widehat{\Delta}_{i}^{n},$
where $I=\sup\{i:\tau_{i}^{n}\leq T\}$. For any $\delta>0$ the event $B_{\delta}^{n}=\{I\in((T-\delta)m_{n},(T+\delta)m_{n})\}$ has probability tending to 1, and on this event we have
$|\rho^{n}(\widetilde{\lambda}^{n}(T))-\widetilde{\sigma}^{n}(T)|\leq\frac{1}{m_{n}}\sum_{i=\lfloor(T-\delta)m_{n}\rfloor}^{\lfloor(T+\delta)m_{n}\rfloor}\widehat{\Delta}_{i}^{n}.$
As $n\to\infty$ the right-hand side converges in distribution to $\widetilde{\sigma}_{T+\delta}-\widetilde{\sigma}_{T-\delta}$. Thus, by picking $\delta$ suitably for each $\varepsilon$, we see that $\mathbb{P}\left(|\rho^{n}(\widetilde{\lambda}^{n}(T))-\widetilde{\sigma}^{n}(T)|>\varepsilon\right)\to 0$. It follows that $d_{D}(\widetilde{\sigma}^{n},\rho^{n})\overset{p}{\to}0$. We also know $d_{D}(\rho^{n},\rho)\to 0$ almost surely, hence in probability and so $\widetilde{\sigma}^{n}\overset{p}{\to}\rho$ in the Skorokhod metric.
We also know $\widetilde{\sigma}^{n}\overset{d}{\to}\widetilde{\sigma}$, so we identify $\widetilde{\sigma}_{t}\overset{d}{=}\sum_{i=1}^{\infty}\Delta_{i}\mathbbm{1}\left(E_{i}\leq\Delta_{i}t\right)$. It remains only to identify $(\Delta_{i})$. By 5.1 we know that for any $k$ we have $a_{n}^{-1}(\Delta_{1}^{n},\dots,\Delta_{k}^{n})\overset{d}{\to}(\Delta_{1},\dots,\Delta_{k})$, so we will show the same sequence converges to $(\Theta_{1},\dots,\Theta_{k})$. Let $L^{n}:\{0,1,\dots,n\}\to\mathbb{Z}$ be the Łukasiewicz path of the discrete tree $T^{n}$ of size $n$, and construct a càdlàg function $e^{n}\in D([0,1])$ by taking
@@
Then we have $e^{n}\overset{d}{\to}\mathbbm{e}$, where $\mathbbm{e}$ is a normalised $\alpha$-stable excursion (this result is implicit in [24] or [33]). Now consider a map $\mathfrak{J}:D([0,1])\to\mathbb{R}^{k}$ given as follows: for any $f\in D([0,1])$, we identify the $k$ largest upward jumps of $f$ and arrange them into decreasing order (appending zeros if there are fewer than $k$ such jumps). This gives a well-defined continuous map, and applying it to the above convergence gives
@@
Appendix A Appendix: omitted proofs
@@
First observe that $\sigma$ is killed at rate $a$, independently of its other jumps, and so $\mathbb{P}\left(\sigma_{t}<\infty\right)=e^{-at}$. We can pull this factor out of the expectation and henceforth forget about the killing in $\sigma$: we aim to show that
@@
To this end, set $G(z)=icz-\Psi_{L}(z)$, $z\in\mathbb{C}$. We can show that, for any closed polygonal contour $\gamma$, we have $\oint_{\gamma}G(z)\,dz=0$ — this is by Fubini’s theorem, which is valid due to the condition that $\int_{0}^{\infty}(x\wedge x^{2})\,\nu(dx)<\infty$. Thus by Morera’s theorem $G$ is holomorphic on $\mathbb{C}$, so $\exp(G(z))$ is holomorphic too.
@@
To this end, we will show that the integrands of these two integrals tend to 0 uniformly: that is, for $z=\lambda-is$, $s\in[0,t]$ we have $\exp(G(z))\to 0$ uniformly in $s$ as $|\lambda|\to\infty$. As $|\exp(G(z))|=\exp(\textnormal{Re}\left(G(z)\right))$ it suffices to show $\textnormal{Re}\left(G(z)\right)\to$-$\infty$ uniformly, and since $s$ is bounded we can further reduce this to showing that $\textnormal{Re}\left(\Psi_{L}(z)\right)\to+\infty$ uniformly. It is straightforward to calculate
@@
$\mathbb{E}\left[\left|\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}T\right)+\mathbbm{1}\left(E_{i}\leq x_{i}T\right)\right)\right|\right]$
$\qquad\qquad\leq\sum_{i=N+1}^{\infty}x_{i}\left(\mathbb{P}\left(E_{i}\leq x_{i}^{(n)}T\right)+\mathbb{P}\left(E_{i}\leq x_{i}T\right)\right)$
$\qquad\qquad\leq T\sum_{i=N+1}^{\infty}x_{i}(x_{i}^{(n)}+x_{i})$
@@
which converges to 0 almost surely (on the event that $\lambda^{(n)}$ is defined). Letting $A_{3}^{(n)}$ be the event that $\sup_{t\in[0,T]}\left|\lambda^{(n)}(t)-t\right|\leq\varepsilon$, we see that $\mathbb{P}\left(A_{3}^{(n)}\right)\to 1$. Note that $\widetilde{R}^{(n)}(\lambda^{(n)}(t))=\widetilde{S}(t)$ for all $t\in[0,T]$. Now we can compute
$\sup_{t\in[0,T]}|R^{(n)}(\lambda^{(n)}(t))-S(t)|$
$\qquad\qquad=\sup_{t\in[0,T]}\left|\sum_{i=1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}\lambda^{(n)}(t)\right)-\mathbbm{1}\left(E_{i}\leq x_{i}t\right)\right)\right|$
$\qquad\qquad=\sup_{t\in[0,T]}\left|\widetilde{R}^{(n)}(\lambda^{(n)}(t))-\widetilde{S}(t)+\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}\lambda^{(n)}(t)\right)-\mathbbm{1}\left(E_{i}\leq x_{i}t\right)\right)\right|$
$\qquad\qquad=\sup_{t\in[0,T]}\left|\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}\lambda^{(n)}(t)\right)-\mathbbm{1}\left(E_{i}\leq x_{i}t\right)\right)\right|$
$\qquad\qquad\leq\left|\sum_{i=N+1}^{\infty}x_{i}\left(\mathbbm{1}\left(E_{i}\leq x_{i}^{(n)}T\right)+\mathbbm{1}\left(E_{i}\leq x_{i}T\right)\right)\right|.$
@@
This completes the proof in the case that $\mathbf{X}^{(n)}$ and $\mathbf{X}$ are deterministic. For the general case, we use Skorokhod’s representation theorem again, this time to construct $\mathbf{X}^{(n)}$ and $\mathbf{X}$ on the same space such that $\mathbf{X}^{(n)}\to\mathbf{X}$ almost surely. Constructing our exponential random variables also on this space, independent of everything else, and using them to define all processes $S^{(n)}$ and $S$, we deduce from the previous case that for any fixed $\varepsilon>0$ we have
@@
Acknowledgements
We are grateful to Andreas Kyprianou for a very helpful suggestion in the early stages of the project, and to Minmin Wang for a discussion of [41]. LH’s work is supported by the EPSRC Centre for Doctoral Training in the Mathematics of Random Systems: Analysis, Modelling, and Simulation (EP/S023925/1). Some of the work was carried out while CG was in residence at the Simons Laufer Mathematical Sciences Institute in Berkeley, California, during the Spring 2025 semester, which LH also visited for a week with funding from the Institute. They gratefully acknowledge the financial support of the National Science Foundation under Grant No. DMS-1928930.